What is the circuit (amps) for a series circuit containing a 12 V potential and three resistors R1= 3 ohms,

R2= 2 ohms and R3= 1 ohm.


1
Expert's answer
2021-05-10T09:17:21-0400

Gives

R1=3 "\Omega"

R2=2 "\Omega"

R3=1"\Omega"

E=12 V

Now

"R_{net}=R_{1}+R_{2}+R_{3}" "R_{net}=(3+2+1) \Omega"

"R_{net}=6\Omega"

"I_{net}=\frac{V}{R_{net}}\rightarrow(1)"

"I_{net}=\frac{12}{6} =2 A"

"I_{net}=2 A"